OpenStudy (anonymous):

Find the correct answer to the nearest tenth 1. 118% of 621 is what number? 2. 0.6% of 180 3. 20 is 125% of what number.

OpenStudy (anonymous):

732,78

myininaya (myininaya):

1) 118/100*621=732.78 to nearest tenth it is 732.8 2) .6/100*180=1.08 to nearest tenth it is 1.1 3) 20=125%n 20=125/100*n 20*100/125=n n=16 to nearest tenth you can leave as 16 or if you like you can write as 16.0